A manager of Waterworks Inc., was reviewing the water bills of a dog daycare and spa. He determined that its highest and lowest bills of $3,800 and $2,400 were incurred in the months of May and November,
respectively.

  
If 400 dogs were washed in May and 200 dogs were washed in November, what was the estimated cost, If 300 dogs were washed in December?

$3,357

$3,573

$3,735

$3,100
Answer #1

Variable cost per dog

= (3800-2400)/(400-200) = 1400/200 = 7 per dog

Fixed cost = Total cost - Variable cost

= 3800 - (400*7) = 3800 - 2800 = 1000 per month

Total Cost for 300 dogs

= (300*7) + 1000

= 2100 +1000

= 3100